# Converging to Stability in Two-Sided Bandits: The Case of Unknown Preferences on Both Sides of a Matching Market

### v2

@@ 4.1. Proof of OCA-UCB Convergence
First, let us define some terminology. A triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ is inconsistent if $p_{i}$ believes $p_{i}\succ_{a_{k}}p_{j}$ but the opposite is true. Then, let us define a plausible set constructed by $p_{i}$ to be inconsistent with respect to $p_{j}$ and $a_{k}$ if $a_{k}\in S_{i}(t)$, denoted by $\hat{S}_{ij}^{k}(t)$. Finally, let us define an inconsistent plausible set as being inconsequential at time $t$ if $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{i}^{m}:a_{m}\in S_{i}(t)\}\neq k$ i.e. $p_{i}$ does not attempt $a_{k}$ OR if $\mathop{\mathrm{arg\,max}}\limits_{m}\{\text{UCB}_{j}^{m}:a_{m}\in S_{j}(t)\}\neq k$ i.e. $p_{j}$ does not attempt $a_{k}$. We begin the proof of the convergence guarantee by first presenting two lemmas:
Lemma 4.0.
Proof.
Lemma 4.0.
Proof.
First, begin by noting that there are only finitely many inconsistent triplets in any given APKP model. Then, a simple recursive argument will suffice to show that each one gets resolved or becomes inconsequential. The first and second cases from Lemma 4.1 are base cases. The first case corrects $p_{i}$ belief altogether, and in the second case the incorrect belief never gets used in any significant way. In the first sub-case of case three, if $a_{k}$ is in-fact $p_{i}$’s best possible achievable arm, then it will continue getting matched with $a_{k}$ regardless of it’s wrong belief (inconsequential w.r.t $a_{k}$). Nonetheless, if $\exists p_{j}$ at $(t+1)$ such that triplet $Q_{ij}^{k}=(p_{i},p_{j},a_{k})$ exists, then we can apply Lemma 4.1 recursively. And finally, in the second sub-case of case three from Lemma 4.1, $p_{i}$ loses the conflict to some other player which implies that some belief gets updated resulting in the resolution of some triplet. However, if $\exists p_{j}$ at $(t+1)$ such that triplet $Q_{i,j}^{k}=(p_{i},p_{j},a_{k})$ exists, then again we can apply Lemma 4.1 recursively. ∎
The proof of Theorem 4.1 is the immediate consequence of lemmas 4.1, and 4.2. Once all the inconsistent plausible sets are either resolved, or are inconsequential, the best arm picked from each player’s plausible set will be the same as the one picked from CA-UCB’s plausible set. This implies that this approach to updaing beliefs has no effect on CA-UCB algorithm’s guarantee on convergence. ∎
We present empirical evidence on the performance of OCA-UCB in Section 8.1.
5. APU Model and PCA-SCA
The previous section looked at the APKP model where the agents did not have access to arms’ preferences. Instead, the players relied on the fact that the arms had perfect knowledge about their own preferences and used deterministic feedback obtained to recreate something similar to the plausible set from the CA-UCB algorithm. However, in the APU model, even arms lack knowledge of their preferences, making feedback received from them noisy, invalidating the approach from the previous section.
Given the issue of unreliable feedback associated with conflict results, we must formulate an approach in which agents can pick the best possible arms while avoiding conflicts. To this end, we use the notion of highest expected reward to determine agent proposals. Each player tracks its estimated probability of winning a conflict against other players. Players choose the arm that maximizes the product of reward estimates with conflict win probability estimates.
@@ 5. APU Model and PCA-SCA
Concurrently, in our algorithm, each arm maintains an empirical reward estimate for each player along with a confidence interval. These confidence intervals serve two purposes: first, when the confidence intervals for a given set of players are disjoint, the arm can deterministically identify the player with the highest expected reward. Second, when the intervals overlap, the arm resolves the conflict by selecting a player uniformly at random among those with overlapping intervals. This allows us to maintain the convergence guarantees of our approach. We detail the algorithmic implementation of this approach in Algorithm 2.
1:Input: $\lambda$, $\kappa$, $H$ 2:Output: A stable match 3:for $t\in\{1,\cdots,H\}$ do 4:     for $i\in\{1,\cdots,N\}$ do 5:         if $t==1$ then 6:              $A_{i}(t)\leftarrow a_{k},k\sim\text{Uniform}(\left[K\right])$ 7:         else 8:              $D_{i}(t)\sim Ber(\lambda)$ 9:              if $D_{i}(t)==0$ then 10:                  $A_{i}(t)\leftarrow$ GetBestArm($p_{i},t$) 11:              else 12:                  $A_{i}(t)\leftarrow A_{i}(t-1)$ 13:              end if 14:         end if 15:     end for 16:     for $k\in\{1,\cdots,K\}$ do 17:         $p_{w}\leftarrow$ ResolveRequests($a_{k},t$) 18:         if $p_{w}\neq\emptyset$ then 19:              $r_{p_{w}},r_{a_{k}}\leftarrow$ Sample-Rewards($p_{w},a_{k}$) 20:              Arm-Update-Belief($a_{k},p_{w},r_{a_{k}}$) 21:              Player-Update-Belief($p_{w},a_{k},r_{p_{w}}$) 22:              $L\leftarrow\left\{p_{j}:A_{j}(t)=a_{k}\right\}\setminus\left\{p_{w}\right\}$ 23:              for $p_{l}\in L$ do 24:                  Update-Win-Probability($p_{l},p_{w},a_{k}$) 25:              end for 26:         end if 27:     end for 28:end for 1:procedure GetBestArm ($p_{i}$, $t$) 2:    for $k\in\{1,\cdots,K\}$ do 3:        $p_{\bar{p}}\leftarrow$ $p_{j}\text{ s.t. }\bar{A}_{j}(t-1)=a_{k}$ 4:        REWARDS$[k]\leftarrow\mathcal{R}_{i}^{k}$triangleright Reward estimate 5:        $P[k]\leftarrow f\left(\hat{P}_{i\bar{p}}^{k}\right)$triangleright Optimism Function 6:    end for 7:    $j\leftarrow\arg\max\left(\text{REWARDS}\circ\text{P}\right)$ 8:    return $a_{j}$ 9:end procedure 1:procedure ResolvePullRequests($a_{k},t$) 2:    $A_{rq}\leftarrow\left\{p_{i}:A_{i}(t)=a_{k}\right\}$$\triangleright$ pull requests 3:    $O_{inf}\leftarrow\emptyset$triangleright players with $\infty$ reward estimate 4:    $O_{bst}\leftarrow\emptyset$triangleright players with best overlapping reward estimate 5:    $I_{L},I_{U}\leftarrow\infty,-\infty$triangleright Keep track of best reward interval 6:    for $p_{i}\in A_{rq}$ do 7:        $\mathcal{R}_{i}^{U}\leftarrow{\tiny\text{Upper bound on reward estimate for }p_{i}}$ 8:        $\mathcal{R}_{i}^{L}\leftarrow{\tiny\text{Lower bound on reward estimate for }p_{i}}$ 9:        if $\mathcal{R}_{i}^{U}=\infty$ then 10:           $O_{inf}=O_{inf}\cup\{p_{i}\}$ 11:           Continue 12:        end if 13:        if $\mathcal{R}_{i}^{L}>I_{U}$ then $\triangleright$ Better interval 14:           $I_{H}\leftarrow\mathcal{R}_{i}^{U}$, $I_{L}\leftarrow\mathcal{R}_{i}^{L}$ 15:           $O_{bst}=\{p_{i}\}$ 16:        else$(I_{U}\leq\mathcal{R}_{i}^{L})\wedge(\mathcal{R}_{i}^{U}\leq I_{L})$triangleright Overlap 17:           $O_{bst}=O_{bst}\cup\{p_{i}\}$ 18:        end if 19:    end for 20:    if $O_{inf}\neq\emptyset$ then 21:        return $\sim\text{Uniform}(O_{inf})$ 22:    else 23:        return $\sim\text{Uniform}(O_{bst})$ 24:    end if 25:end procedure
The algorithm is parameterized by $\lambda\in[0,1)$, and $\kappa\geq 1$. The first parameter $\lambda$ is used to introduce a random delay mechanism to reduce the likelihood of conflicts (Liu et al., 2021; Kong et al., 2022). This is controlled in line 8 of PCA-SCA where each player draws a Bernoulli Random Variable $D_{i}(t)$ with expectation $\lambda$. If $D_{i}(t)=0$, the player attempts the arm with the highest expected reward, else it will attempt the arm that it did in the previous time step. The second parameter $\kappa$ is used to control the level of optimism that the player uses when determining the probability of it winning a conflict against another player for an arm. This is used in line 5 of procedure GetBestArm. We will shortly detail the specifics of how this level of optimism is determined. Initially all reward estimates are set to infinity, and all win conflict probabilities are set to 1. First let us go over how the arms pick the players at each round.
@@
Each arm processes the incoming proposals as follows. If at least one player’s reward estimate is still infinite (i.e., uninitialized), the arm selects uniformly at random from those players (Line 21 of ResolvePullRequests). Otherwise, the arm identifies the player with the highest upper confidence bound and checks whether other players’ confidence intervals overlap with this best bound (Line 16). If multiple intervals overlap, the arm selects uniformly at random from among those candidates (Line 23); if only one player remains with the highest bound, it selects that player.
@@
As in CA-UCB, each player begins with all arm reward estimates set to $\infty$. However, players also maintain probability estimates $\hat{P}_{ij}^{k}$, which represent player $p_{i}$’s belief of winning a conflict against player $p_{j}$ for arm $a_{k}$. These estimates are updated based on the ratio of conflicts won to total conflicts encountered and are initialized to 1.
Over time, players gather conflict outcomes and adjust $\hat{P}_{ij}^{k}$ accordingly. At each time step, $p_{i}$ selects the arm that maximizes its expected reward:
$\arg\max_{k}\Bigl{[}\hat{P}_{ij}^{k}\cdot\mathcal{R}_{i}^{k}\Bigr{]},$
where $j$ is the index of the player who pulled arm $k$ in the previous step, and $\mathcal{R}_{i}^{k}$ is $p_{i}$’s current reward estimate for arm $a_{k}$.
Remember that while the arms’ confidence intervals overlap, they select players uniformly at random, rendering conflict feedback noisy. In expectation, $\hat{P}_{ij}^{k}$ converges to $0.5$ during this phase. Once the intervals become disjoint, selection becomes deterministic, allowing $\hat{P}_{ij}^{k}$ to approach its true value. However, a player with an “unlucky” estimate that drifts below $0.5$ may prematurely abandon that arm. To encourage continued exploration—even under unfavorable conflict estimates—we introduce the optimism function:
(1) $f(x)=\begin{cases}\displaystyle\frac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&0\leq x\leq 0.5,\\[4.30554pt] 1,&x>0.5,\end{cases}$
where $\kappa\geq 1$ is a hyper-parameter that controls the level of optimism in the algorithm. Equation 1 is employed to modulate optimism due to its favorable behavior with respect to each $\hat{P}_{ij}^{k}$. Specifically, if $\hat{P}_{ij}^{k}>0.5$, the player is fully optimistic about its win chances; however, when $\hat{P}_{ij}^{k}$ falls below 0.5, the function $f(\hat{P}_{ij}^{k})$ boosts the estimate to 1 and then decreases it gradually, with the rate of decrease governed by $\kappa$. This behavior is illustrated in Figure 6 (in the Appendix), which plots $f(x)$ for different values of $\kappa$.
As $\kappa\to\infty$, $f(\hat{P}_{ij}^{k})$ approaches 1 regardless of the value of $\hat{P}_{ij}^{k}$, whereas $\kappa=1$ yields an almost linear, yet optimistic, dropoff. As players begin receiving deterministic feedback, $\hat{P}_{ij}^{k}$ converges to the true win probability; if that value is 1, then any estimate with $\hat{P}_{ij}^{k}\geq 0.5$ is boosted to 1 by the function, while if the true probability is 0, the function decreases the estimate slowly—again, at a rate determined by $\kappa$. Thus, the choice of $\kappa$ should reflect how aggressively one wishes to maintain optimism.
With this information, at each subsequent time step, the players will attempt to pull the arm that maximizes the product of the reward estimate and the optimistic estimate of the probability of winning (line 7 of procedure GetBestArm).
Then, at the end of each time step, the arms will pick the player that it believes will give the highest reward (similar to how arms behave in the simultaneous choice mechanism of (Das and Kamenica, 2005)) – with the added caveats mentioned above. A player and arm both receive a reward and update their beliefs when an arm is pulled successfully. No reward is given if an arm is not pulled or if a pull request is not received. Players update their beliefs about winning after all matches are made.
@@
Finally, the structure of our algorithm accommodates multiple methods for updating beliefs about expected rewards and conflict win probabilities. In this work, we consider two approaches: UCB, as employed in (Liu et al., 2021), and a Thompson Sampling variant, as presented in (Kong et al., 2022). We prove that using UCB for belief maintenance ensures convergence to stability and support this claim with simulation results. We present simulation results for the Thompson Sampling variant, leaving a theoretical convergence proof for future work.
6. Using UCB in PCA-SCA (PCA-UCB)
In this section, we look at using the Upper Confidence Bound (UCB) approach to keep track of reward estimates, as used in the multi-armed bandit literature by (Auer et al., 2002). We use UCB in the PCA-SCA algorithm and refer to it as PCA-UCB. Players must keep track of beliefs about payoffs and the probability of winning conflicts. Equation 2 describes how beliefs about reward payoffs are estimated by the players.
@@ 6. Using UCB in PCA-SCA (PCA-UCB)
(2) $\text{UCB}_{i}^{k}(t)=\begin{cases}\infty&\qquad\text{ if }n_{i}^{k}(t)=0\\ \hat{\mu}_{i}^{k}+\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t-1)}}&\qquad\text{ otherwise}\end{cases}$
where $n_{i}^{k}(t)$ is the number of times player $p_{i}$ has pulled arm $a_{k}$ at time-step $t$, and $\hat{\mu}_{i}^{k}$ is the empirical mean tracked by player $p_{i}$ for arm $a_{k}$. Similarly, the arms keep track of the confidence intervals for the rewards offered by the players. This is done by having the arms keep track of both the UCB and the Lower Confidence Bound (LCB) associated with each player:
(3) $\begin{array}[]{rcl}\text{LCB}_{i}^{k}(t)&=&\hat{\mu}_{i}^{k}(t)-c_{i}^{k}(t),\\ \text{UCB}_{i}^{k}(t)&=&\hat{\mu}_{i}^{k}(t)+c_{i}^{k}(t)\end{array}$
@@
where $c_{i}^{k}(t)=\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t)}}$. Similar to the players, the LCB and the UCB for the arms is initialized to $\infty$- ∞ and $\infty$ respectively. The other quantity that the players keep track of is the probability of conflict wins:
(4) $\hat{P}_{ij}^{k}(t)=\begin{cases}1&\qquad\text{ if }n_{ij}^{k}=0\\ \frac{w_{ij}^{k}(t)}{N_{c_{ij}}^{k}(t)}&\qquad\text{ otherwise}\end{cases}$
where $\hat{P}_{ij}^{k}(t)$ represents player $p_{i}$’s belief about the probability of winning a conflict against player $p_{j}$ for arm $a_{k}$ at time $t$. $w_{ij}^{k}$ is the number of times player $p_{i}$ has won this conflict and $N_{c_{ij}}^{k}(t)$ is the total number of times this conflict has happened thus far. At the end of each time step, belief estimates are updated once rewards are sampled and matching information is made public. If a player wins a conflict, the corresponding probability estimate increases, otherwise, it decreases. This is handled by the Update-Win-Probability function in Algorithm 2.
@@
Before presenting the sketch of the proof of convergence, we will first provide some preliminary definitions and notation. First, in our analysis, we assume without loss of generality that players are ordered so that $\mu_{i}^{k}>\mu_{j}^{k}$ when $p_{i}\succ_{a_{k}}p_{j}$. We say that the confidence intervals for arm $a_{k}$ are disjoint if, for every pair $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$,
$\text{LCB}_{i}^{k}(t)>\text{UCB}_{j}^{k}(t).$
Conversely, if for some $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$ we have
$\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t),$
then the confidence intervals for those two players are said to be overlapping. As mentioned before, when an arm’s confidence intervals overlap for a set of players, the conflict feedback is provided uniformly at random. Specifically, the arm selects the proposal with the highest UCB estimate; if there is any overlap with another player’s interval, it chooses uniformly at random from the contenders. Since pairwise win probability estimates are only updated when one of the two members of the pair actually wins the arm at that time (thus conflicts between more than two players do not confound the estimate), the win probability estimates for losing players concentrate around $0.5$. Once the intervals become disjoint, however, the feedback becomes deterministic. Let us now define this deterministic feedback event:
@@
In CA-UCB, players had direct access to arms’ preferences, whereas in OCA-UCB, the arms’ preferences were learned deterministically. In the context of the APKP model, one may view an arm $a_{k}\in S_{i}(t)$ as having a win probability of 1 (with all other arms having a win probability of 0), thereby considering arms with a win probability of 1 as “plausible”. In the APU model, however, plausibility must be interpreted more generally as any arm with a non-zero probability of being won.
@@
[$\Bigl{[}\Bigl{\{}\Bigl{(}P_{ij}^{k}=1\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\leq 1-\epsilon\Bigr{)}\Bigr{\}}\vee\Bigl{\{}\Bigl{(}P_{ij}^{k}=0\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\geq\epsilon\Bigr{)}\Bigr{\}}\Bigr{]},$
i.e., the player’s empirical belief about its win probability in the conflict deviates significantly (by at least $\epsilon$) from the true outcome—either overestimating a win or underestimating a loss.
Next, we define a inconsequential triplet
@@
A triplet is considered consequential if it is not inconsequential. And finally, for any pair of players $p_{i}$, $p_{j}$, and $a_{k}$, define:
@@
Note that, without loss of generality, in our models, we assume $\Delta_{\min}=1$. With these preliminaries out of the way, we will now proceed to provide a sketch of the proof that proves that using UCB in the PCA-SCA algorithm yields convergence to market stability. A full proof can be found in Appendix A.
6.1. Sketch of the Proof of Convergence to Stability Using PCA-UCB
Consider an APU model with $K$ arms and $N$ players. Our approach is to show that (i) each arm learns its true preference ordering over players in finite time, (ii) once the arms’ estimates are sufficiently accurate, the conflict feedback becomes deterministic so that the players’ empirical win probability estimates (that is, the estimated probability of winning a conflict against a specific other player for a specific arm) converge to their true values (either 0 or 1), and (iii) as a consequence, the execution of our PCA-UCB algorithm becomes identical to that of CA-UCB in the long run.
The first step is to establish that each arm maintains an empirical reward estimate for every player together with a confidence interval that shrinks as the number of samples increases. By bounding the probability that the confidence intervals of two players overlap—that is, that the lower confidence bound of a higher-mean player exceeds the upper confidence bound of a lower-mean player—we show that the probability of any overlap decays exponentially with the number of observations. We denote the event that all arms have learned their preferences by time $t$ as $\alpha(t)$.
Conditioned on the event $\alpha(t)$, the feedback that players receive during conflicts becomes deterministic. In the early stages, when the arms’ confidence intervals may overlap, the feedback is random and the players’ win estimates concentrate around $0.5$. To prevent premature abandonment of an arm due to these fluctuations, our algorithm incorporates an optimism function that boosts the win probability estimates. Once the feedback is deterministic, each conflict yields a noise-free observation, and a straightforward averaging argument shows that the error in the players’ empirical win estimates decreases proportionally to $\frac{T}{T+t}$ after time $T$.
Finally, by combining the bounds on the probability that the arms’ confidence intervals overlap with the convergence rate of the win estimates, we conclude that the overall failure probability vanishes over time. In this asymptotic regime, the decision-making process of PCA-UCB exactly replicates that of CA-UCB. Now, we will provide the sketch of the proof of each of the theorems.
@@ 6.1. Sketch of the Proof of Convergence to Stability Using PCA-UCB
theoremTheoremOne[Finite-Time Learning of Arms’ Preferences] In an instance of the APU model with $K$ arms and $N$ players running PCA-UCB, at time $t$, with probability at least $1-\delta(t)$, for every arm, the confidence intervals that the arm tracks for each player are disjoint. Here $\delta(t)$ is :
$\delta(t)=K\cdot\frac{N^{2}}{2}\cdot\exp\left(-\frac{n(t)\left(\Delta_{\text{min}}-2c(t)\right)^{2}}{4}\right)$
where $n(t)=\underset{\begin{subarray}{c}i\in N\\ k\in K\end{subarray}}{\min}\left\{n_{i}^{k}(t)\right\}$, $\Delta_{\min}=\underset{\begin{subarray}{c}i,j\in N\\ k\in K\end{subarray}}{\min}\left\{\Delta_{ij}^{k}\right\}$, and $c(t)=\sqrt{\frac{3\ln(t)}{2n(t)}}$.
To prove Theorem 6.1, we first analyze the overlap probability for a single pair of players on a fixed arm. By writing the event of overlapping confidence intervals as
$\hat{\mu}_{i}(t)-\hat{\mu}_{j}(t)\leq c_{i}(t)+c_{j}(t),$
and by applying a concentration inequality (specifically, a Hoeffding bound for sub-Gaussian random variables), we obtain an exponentially decaying bound on the probability that the intervals for any pair do not separate. A union bound over all player pairs and over all $K$ arms then shows that the probability that any arm has overlapping intervals is bounded by $\delta(t)$. In other words, as the number of samples $n(t)$ grows, the event $\alpha(t)$—where every arm’s confidence intervals are disjoint—occurs with high probability.
@@
Here $\tau(t)$ is:
$\frac{KN^{2}}{2}\exp\left(-2\,N_{c}(T)\,\left(\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}\right)^{2}\right)$
where $N_{c}(t)$ is the minimum number of conflicts across all triplets.
Theorem 6.1 is proved by first noting that prior to $\alpha$ the feedback provided by the arms is random, leading to win estimates that concentrate around $0.5$. To ensure that players continue to sample even when their estimates are slightly off, we introduce an optimism function $f(x)$ that amplifies estimates near $0.5$. Once $\alpha$ has occurred, the conflict outcomes become deterministic so that each conflict provides a true (noise-free) observation. Using a simple averaging argument, one can then show that the error in the win estimate decays as $\frac{T}{T+t}$. Moreover, any triplet for which the win estimate does not converge sufficiently fast is rendered inconsequential to the player’s decision-making. A careful matching of the tail probabilities of the concentration inequalities yields the bound $\tau(t)$.

### v4

@@ 3 APKP Model and OCA-UCB
4 APU Model and PCA-SCA
The previous section studied the APKP model, where players lacked direct knowledge of arms’ preferences but leveraged deterministic feedback—obtained from arms that perfectly knew their preferences—to reconstruct a plausible set akin to CA-UCB. In contrast, under the APU model, arms themselves do not know their preferences, leading to noisy conflict feedback and invalidating the previous approach. Due to this unreliable conflict feedback, we propose an alternative approach enabling agents to select optimal arms while avoiding conflicts. We introduce a selection rule based on maximizing the expected reward combined with estimated probabilities of winning conflicts. Each player maintains estimates of its conflict-win probabilities against other players, choosing the arm that maximizes the product of these probabilities with reward estimates.
@@ 4 APU Model and PCA-SCA
Simultaneously, each arm maintains empirical reward estimates and confidence intervals for all players. These intervals serve two purposes: when intervals are disjoint, the arm deterministically identifies the highest-reward player; when intervals overlap, the arm resolves conflicts by selecting among overlapping players uniformly at random. This approach ensures convergence. The full algorithm (Algorithm 2) can be found in the Appendix.
The algorithm uses two parameters, $\lambda\in[0,1)$ and $\kappa\geq 1$. $\lambda$ injects randomized delays (line 8 of PCA-SCA) by sampling $D_{i}(t)\sim\mathrm{Bernoulli}(\lambda)\,$ to reduce the likelihood of conflicts Liu et al. (2021); Kong et al. (2022). If $D_{i}(t)=0$, player $i$ selects the arm with the highest expected reward; otherwise, it repeats its previous choice. $\kappa$ tunes the level of optimism in the player’s conflict–win probability estimates (line 5 of GetBestArm) – the specifics of which are detailed below. All reward estimates are initialized to $\infty$ and all win probabilities to 1. First let us go over how the arms pick the players at each round.
@@
Each arm processes the incoming proposals as follows. If at least one player’s reward estimate is still infinite (i.e., uninitialized), the arm selects uniformly at random from those players (Line 30 of ResolvePullRequests). Otherwise, the arm identifies the player with the highest upper confidence bound and checks whether other players’ confidence intervals overlap with this best bound (Line 25). If multiple intervals overlap, the arm selects uniformly at random from among those candidates (Line 32); if only one player remains with the highest bound, it selects that player.
@@
Each player initializes all arm-reward estimates to $\infty$ and all empirical conflict-win probabilities $\hat{P}_{ij}^{k}=1$, updating each $\hat{P}_{ij}^{k}$ to the fraction of conflicts won. At each time step, player $p_{i}$ selects the arm that maximizes its expected reward $\arg\max_{k}\Bigl{[}\hat{P}_{ij}^{k}\cdot\mathcal{R}_{i}^{k}\Bigr{]}$, where $j$ is the index of the player who pulled arm $k$ in the previous step, and $\mathcal{R}_{i}^{k}$ is $p_{i}$’s current reward estimate for arm $a_{k}$. When there is overlap, arms choose uniformly at random, so in expectation $\hat{P}_{ij}^{k}\to 0.5$. Once intervals separate, selection is deterministic and $\hat{P}_{ij}^{k}$ converges to its true value. To prevent premature arm abandonment from transiently low estimates, we introduce the optimism function:
$f(x)=\begin{cases}\dfrac{1-e^{-\kappa x}}{1-e^{-\kappa/2}},&0\leq x\leq 0.5,\\[4.30554pt] 1,&x>0.5\end{cases}$ (1)
Here $\kappa\geq 1$ controls the level of optimism. As $\kappa\to\infty$, $f(\hat{P}_{ij}^{k})\to 1$ for all $\hat{P}_{ij}^{k}$, whereas $\kappa=1$ yields a nearly linear yet optimistic taper. Once feedback is deterministic and $\hat{P}_{ij}^{k}\to p_{ij}^{k}$, any $\hat{P}_{ij}^{k}\geq 0.5$ is boosted to 1 when $p_{ij}^{k}=1$, while for $p_{ij}^{k}=0$, $f(\hat{P}_{ij}^{k})$ decays slowly at a rate determined by $\kappa$. Thus, $\kappa$ directly controls the aggressiveness of optimism (see Figure 5 in Appendix F).
At each round, every player selects the arm that maximizes the product of its reward estimate and its optimistic win-probability (line 7 of GetBestArm). Then each arm chooses the player with the highest expected reward (mirroring the simultaneous-choice rule of Das and Kamenica (2005)). Successful pulls generate mutual rewards and update both arms’ reward estimates and players’ win-probability beliefs.
@@
Finally, note that the structure of our algorithm accommodates multiple methods for updating beliefs about expected rewards and conflict win probabilities. Here, we test UCB Liu et al. (2021) and Thompson Sampling Kong et al. (2022) for belief updates in simulation, in addition to providing a proof of convergence to stability when UCB is used.
4.1 Using UCB in PCA-SCA (PCA-UCB)
We integrate the classic Upper Confidence Bound (UCB) Auer et al. (2002) method into PCA-SCA (see Algorithm 2) and refer to it as PCA-UCB. Each player maintains
@@ 4.1 Using UCB in PCA-SCA (PCA-UCB)
$\text{UCB}_{i}^{k}(t)=\begin{cases}\infty&\qquad\text{ if }n_{i}^{k}(t)=0\\ \hat{\mu}_{i}^{k}+\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t-1)}}&\qquad\text{ otherwise}\end{cases}$ (2)
@@
where $n_{i}^{k}(t)$ is the number of times player $p_{i}$ has pulled arm $a_{k}$ at time-step $t$, and $\hat{\mu}_{i}^{k}$ is the empirical mean tracked by player $p_{i}$ for arm $a_{k}$. Let $c_{i}^{k}(t)=\sqrt{\frac{3\ln(t)}{2n_{i}^{k}(t-1)}}$, then each arm tracks for every player both a lower and upper confidence bounds given by $\text{LCB}_{i}^{k}(t)=\hat{\mu}_{i}^{k}(t)-c_{i}^{k}(t)$ and $\text{UCB}_{i}^{k}(t)=\hat{\mu}_{i}^{k}(t)+c_{i}^{k}(t)$. Similar to the players, the LCB and the UCB for the arms is initialized to -$\infty$ and $\infty$ respectively. The other quantity that the players keep track of is the probability of conflict wins:
$\hat{P}_{ij}^{k}(t)=\begin{cases}1&\qquad\text{ if }n_{ij}^{k}=0\\ \frac{w_{ij}^{k}(t)}{N_{c_{ij}}^{k}(t)}&\qquad\text{ otherwise}\end{cases}$ (3)
Here, $\hat{P}_{ij}^{k}(t)$ is $p_{i}$’s estimated probability of defeating $p_{j}$ on arm $a_{k}$, where $w_{ij}^{k}(t)$ and $N_{c_{ij}}^{k}(t)$ are the counts of wins and total conflicts. After each round—once rewards and matchings are revealed—the Update-Win-Probability function (Algorithm 2) increments the estimate on a win and decrements it on a loss.
@@
To state our convergence theorems, we introduce the following notation and definitions. We assume, without loss of generality, that players are ordered so that $\mu_{i}^{k}>\mu_{j}^{k}$ when $p_{i}\succ_{a_{k}}p_{j}$. We say that the confidence intervals for arm $a_{k}$ are disjoint if, for every pair $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$, $\text{LCB}_{i}^{k}(t)>\text{UCB}_{j}^{k}(t)$. Conversely, if for some $\{p_{i},p_{j}\}$ with $\mu_{i}^{k}>\mu_{j}^{k}$ we have $\text{LCB}_{i}^{k}(t)\leq\text{UCB}_{j}^{k}(t),$ then the confidence intervals for those two players are said to be overlapping. Next, the deterministic feedback event is defined as:
@@
CA-UCB assumes known preferences, and OCA-UCB learns them deterministically—treating arms in $S_{i}(t)$ as having win probability of 1. In APU, plausibility generalizes to any arm with positive win probability. Therefore, we define inconsistency in the APU model as follows.
@@
Let $P_{ij}^{k}$ denote the true probability of player $p_{i}$ winning against player $p_{j}$ in a conflict for arm $a_{k}$, and let $\hat{P}_{ij}^{k}$ be the corresponding empirical probability. The triplet $(p_{i},p_{j},a_{k})$ is said to be inconsistent if
$\Bigl{[}\Bigl{\{}\Bigl{(}P_{ij}^{k}=1\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\leq 1-\epsilon\Bigr{)}\Bigr{\}}\vee\Bigl{\{}\Bigl{(}P_{ij}^{k}=0\Bigr{)}\wedge\Bigl{(}\hat{P}_{ij}^{k}\geq\epsilon\Bigr{)}\Bigr{\}}\Bigr{]},$
@@
A triplet is considered consequential if it is not inconsequential. Finally, for any pair of players $p_{i}$, $p_{j}$, and $a_{k}$, define:
@@
Note that, without loss of generality, we assume $\Delta_{\min}=1$. With the key definitions in place, we can now state our three main convergence theorems. Their proofs are deferred to the Appendix.
@@
theoremTheoremOne[Finite-Time Learning of Arms’ Preferences] In an instance of the APU model with $K$ arms and $N$ players running PCA-UCB, at time $t$, with probability at least $1-\delta(t)$, for every arm, the confidence intervals that the arm tracks for each player are disjoint. Here $\delta(t)=K\cdot\frac{N^{2}}{2}\cdot\exp\left(-\frac{n(t)\left(\Delta_{\text{min}}-2c(t)\right)^{2}}{4}\right)$ where $n(t)=\underset{\begin{subarray}{c}i\in N\\ k\in K\end{subarray}}{\min}\left\{n_{i}^{k}(t)\right\}$, $\Delta_{\min}=\underset{\begin{subarray}{c}i,j\in N\\ k\in K\end{subarray}}{\min}\left\{\Delta_{ij}^{k}\right\}$, and $c(t)=\sqrt{\frac{3\ln(t)}{2n(t)}}$.
Our first result shows that arms rapidly learn their true preference rankings, so that after a finite number of samples, conflicts between players send deterministic feedback.
@@
Here $\tau(t)=\frac{KN^{2}}{2}\exp\left(-2\,N_{c}(T)\,\left(\left(\Delta_{\text{min}}-2c(T)\right)\sqrt{\frac{n(T)}{8N_{c}(T)}}\right)^{2}\right)$, where $N_{c}(t)$ is the minimum number of conflicts across all triplets.
Theorem 3 guarantees that once feedback becomes noise-free, players’ empirical win-probabilities converge to their true values at a $O(T/(T+t))$ rate.